# Region-aware Grasp Framework with Normalized Grasp Space for 6-DoF Grasping in Cluttered Scene

### v1

@@ Abstract
Regional geometric information is crucial for determining grasp poses. A series of region-based methods succeed in extracting regional features and enhancing grasp detection quality. However, faced with a cluttered scene with multiple objects and potential collision, the definition of the grasp-relevant region remains inconsistent among methods, and the relationship between grasps and regional spaces remains incompletely investigated. In this paper, from a novel region-aware and grasp-centric viewpoint, we propose Normalized Grasp Space (NGS), unifying the grasp representation within a normalized regional space. The relationship among the grasp widths, region scales, and gripper sizes is considered and empowers our method to generalize to grippers and scenes with different scales. Leveraging the characteristics of the NGS, we find that 2D CNNs are surprisingly underestimated for complicated 6-DoF grasp detection tasks in clutter scenes and build a highly efficient Region-aware Normalized Grasp Network (RNGNet). Experiments conducted on the public benchmark show that our method achieves the best grasp detection results compared to the previous state-of-the-arts while attaining a real-time inference speed of approximately 50 FPS. Real-world cluttered scene clearance experiments underscore the effectiveness of our method with a higher success rate than other methods. Further human-to-robot handover and moving object grasping experiments demonstrate the potential of our proposed method for closed-loop grasping in dynamic scenarios.
I Introduction
@@
Grasping is a fundamental task for robotic manipulation and many downstream applications. Recent advancements in deep-learning methodologies enable data-driven grasp detection techniques to directly deduce the positions and rotations of available grasps, achieving better generalization to novel scenes compared with the classical template-based grasp detection methods. Meanwhile, 6-DoF grasping has grown significantly because of its effectiveness in applications and scenarios that require accurate and reasonable manipulation of objects.
@@
Towards efficient 6-DoF grasping, most grasp detection methods treat the task as a detection-style problem, involving the acquisition of optimal grasp positions and spatial rotations from a single-view RGBD image or point cloud. Most of them process the global scene information to generate all possible grasps in the scene. Typical works like [30, 35] extract scene geometric features and regress grasps directly. The crucial problem is the redundancy and ambiguity of the entire scene information, which is time-consuming and demonstrates limited efficacy. Facing the challenge of grasping in cluttered scenes, recent methods [9, 38, 46, 40, 21, 24] encode global per-points features to infer potential grasp centers and segment graspable regions by integrating the regional points with the global features. While global-to-local schemes have demonstrated effectiveness, leveraging regional information without a consistent representation may undermine the generalization capability for novel scenes. The primary challenge lies in devising a method to extract regional information in a consistent representation.
Figure 1: We propose a Region-aware Grasp Framework for 6-DoF grasping in cluttered scenes by detecting regional grasps within RGBD patches. The Normalized Grasp Space (NGS) is a 3D space normalized by the given gripper width, represented as an adaptively sampled and normalized image patch. Based on the depth of the patch center, we adaptively extract regional patches from the original RGBD image. These normalized patches are then fed to the Regional Normalized Grasp Network (RNGNet) to infer regional grasps.
@@
As is shown in Fig. 1, in this work, we propose a novel Region-aware Grasp Framework decoupling grasp detection into graspable region extraction and follow-up regional grasp generation. Specifically, we propose Normalized Grasp Space (NGS), a unified representation bridging the patch extraction and regional grasp detection. To fit the new grasp representation, we reconstruct a regional grasping dataset from the widely adopted GraspNet-1Billion dataset [9], consisting of numerous region patches and grasp annotations. With extracted region information, the grasp detection problem within NGS can be formulated into a normalized form, which shows the invariance to the gripper size and the robustness to perturbations on the grasp center localization. Thus, by considering the grasp detection from a region-aware grasp-centric viewpoint, we further propose Regional Normalized Grasp Network (RNGNet) to extract regional grasp-relevent geometric features and generate regional grasps, which demonstrates enhanced robustness and efficiency. Quantitative experiment results show a significant over 20% improvement in the grasp Average Precision (AP) on the widely used public benchmark with a lower computation cost and faster inference speed than the former SOTA, AnyGrasp [10]. Real-world experiments including clutter clearance and multi-scale grasp inference, validate our framework’s robustness and generalization capability. Further demonstration of grasp detection in dynamic scenes proves the potential and flexibility of our framework for more complicated tasks.
The contributions of the paper mainly include:
  1. 1)
@@
    We propose a novel Region-aware Grasp Framework for 6-DoF grasp detection in cluttered scenes, decoupling grasp detection tasks and enhancing grasp detection efficiency and generalizability.
  2. 2)
@@
    Normalized Grasp Space (NGS), a novel unified space defined in the camera frame for general parallel gripper grasp detection, empowering aligned and scale-invariant grasp detection in local regions.
  3. 3)
@@
    A highly efficient patch-based Region Normalized Grasp Network (RNGNet) aiming at extracting region-aware features and predicting high-quality 6-DoF grasps inside the NGS.
II Related Work
II-A Vision-based Grasp Detection
By the degree of freedom at the robot gripper, grasp detection methods can be categorized into two main types: 4-DoF grasp detection based on image input, and 6-DoF grasp detection, which predominantly utilizes 3D geometry input. 4-DoF methods encode image features directly and predict grasps as planar rectangles within the image plane [26, 28, 42]. Though efficient, limited by the grasp approach direction and lack of 3D geometry analysis, 4-DoF methods often struggle with complex objects or cluttered scenes and are prone to collision.
@@ II-A Vision-based Grasp Detection
In recent years, another mainstream line of research has focused on detecting 6-DoF grasps to address more challenging scenarios, primarily relying on single-view RGBD images or point clouds. Pioneer works like ten Pas et al. [39] and Liang et al. [18] mainly follow a sample-evaluation strategy, generating numerous grasp proposals and selecting optimal grasps through an evaluation model. Subsequently, Qin et al. [35] employ PointNet++ [33] for per-point feature extraction and direct regression of grasp parameters. Diverging from existing approaches, our method utilizes a streamlined Convolutional Neural Network (CNN) architecture to extract regional information directly from RGBD images. This transition from scene-level to region-level analysis simplifies the task, facilitating the generation of high-quality grasps with improved efficiency.
Recent advancements, benefiting from the proposed large-scale datasets [9, 8, 27], have adopted an end-to-end strategy to generate 6-DoF grasps in cluttered scenes. Wang et al. [40] introduce point-wise graspness to represent grasp location probabilities and subsequently aggregate regional features for rotational attributes prediction. HGGD [3] represents the work most closely related to ours, which employs 2D CNNs to generate heatmaps for grasp locations and integrate 2D semantics with 3D geometric features to configure 6D grasps. Our work differs from HGGD in several aspects: 1) We employ a region-aware scheme to detect grasps, forcing our method to capture grasp-relevant and scene-independent regional features. This significantly enhances its generalization capability. However, HGGD relies on scene-level features to deduce grasp attributes. 2) Derived from local regions and considering the physical dimensions, the Normalized Grasp Space exhibits advantageous properties such as translation invariance and rotation equivariance, which are beneficial for grasp detection. Trained from regional data, Region Normalized Grasp Network not only demonstrates improved efficiency compared to HGGD but also readily adapts to more challenging tasks, such as dynamic grasping and grasping object with different scales.
II-B Region-aware Representation for Robotic Manipulation
Recently, considerable research has been dedicated to object descriptors or pose estimators [7, 6, 41]. Although these object-centric representations provide sample-efficiency benefits, these methods require object-specific training data and are faced with difficulty in generalizing to a real-world cluttered scene. In contrast to the ideal environment in simulation, robotic manipulation tasks in real-world scenarios present challenges such as incompleteness, occlusion, and noise, making feature extraction for manipulation considerably more intricate.
Recent research has shown a shift towards a manipulation-centric approach and demonstrating superior generalization capabilities for unseen scenes and objects. For instance, Zeng et al. [44] rearrange deep manipulation features based on regional object and end-effector information to infer actions. Simeonov et al. [37] introduce Neural Descriptor Fields (NDFs) to encode object representations using both points and relative poses, leveraging spatial symmetries of regional representations and showcasing their effectiveness. Recent advancements have also emphasized equivariance. Huang et al. [14] and Zhu et al. [47] extract inherent SE(2) equivariance through 4-DoF planar grasp detection and imitation learning tasks, respectively. Moreover, Simeonov et al. [37] and Ryu et al. [36] design equivariant models to capture SE(3)-equivariance of NDFs in robotic manipulation tasks.
Inspired by the manipulation-centric perspective, we rethink the grasp detection problem from a grasp-centric view. Specifically, we reformulate the problem in a normalized grasp space by focusing on grasps within local regions and propose a novel Region-aware Grasp Framework. The space exhibits favorable equivariance properties, facilitating the framework for accurate and fast inference. Additionally, this configuration offers flexibility for advanced tasks such as dynamic grasping and multi-scale grasp generation.
II-C Region Normalization for Grasp Detection
The utilization of local image patches or regional points is common for both 4-DoF and 6-DoF grasp detection methods [47, 35] recently. Similar to object detection, typical 4-DoF grasp detection approaches [31, 11] conduct image-based normalization utilizing STN [16] to improve model performance and robustness. However, the conventional 4-DoF grasp detection framework is inherently limited, as it primarily focuses on scene-level image normalization, overlooking critical physical factors and 3D geometric information, and thus it might be sensitive to camera distance and viewpoints. In contrast, our approach shifts focus to a 3D regional context instead of the entire scene. We normalize specified regions and their corresponding grasps according to the actual physical sizes of the regions and the gripper. This consistent 3D normalized grasp space allows our proposed Region Normalized Network to excel in general clutter picking tasks and easily adapt to more challenging applications, such as dynamic grasping.
For 6-DoF methods, most methods operate under a fixed gripper setting, extracting local regions of predetermined sizes to detect 6-DoF grasps [9, 40, 46]. However, these methods struggle to generalize or require model retraining when applied to robots with different gripper sizes or to objects of different scales. In contrast, our approach proposes the normalization process to obtain the Regional Grasp Space with beneficial characteristics such as translation and scale invariance. This empowers our framework to generalize to scenarios involving diverse scales.
In addition, relying solely on images for representation frequently lacks comprehensive geometric information, which is available in point clouds or voxels [3, 46]. This limitation may lead to normalization failure due to variations in distance from the scene to the camera or changes in the size of target objects. To address these challenges, we propose a depth-aware technique to extract adaptive patches, allowing for a more coherent patch-based representation for 6-DoF grasping.
@@ II-C Region Normalization for Grasp Detection
Figure 2: Region-aware Grasp Framework Overview. Starting with an RGBD image input, we employ a Grasp Heatmap Model (GHM) [3] to identify potential patch centers. Subsequently, we extract depth-adaptive local patches using an adaptive grid generator, which samples values from the original RGBD image based on the depth of the patch center and the camera’s focal length. These adaptive local patches are then normalized and fed into the Regional Normalized Grasp Network (RNGNet) to detect regional grasps. To obtain grasps back in the camera frame, we apply the inverse of the normalization process. The grasps extracted from all patches collectively form the scene grasps.
III Problem Formulation
@@
In this paper, we focus on the problem of single-view 6-DoF grasp detection of parallel grippers in clutter scenes, in which the camera frame is parallel with the gripper frame. Similar to the problem definition in prior works [3, 4, 9], we aim to generate available grasp poses in the scene with a single-view RGBD image $\boldsymbol{\chi}\in\mathbb{R}^{H\times W\times 4}$ as input. Inspired by [3] and aligned with other definitions, we adopt the grasp representation as:
$\boldsymbol{g}=(x,y,z,\theta,\gamma,\beta,w),$ (1)
in which $(x,y,z)$ is the 3-DoF translations of grasp gripper centers in the camera frame, $w$ is the grasp width and $(\theta,\gamma,\beta)$ compose the 3-DoF intrinsic rotation as Euler angle in the camera frame. $w$ is predicted within the range of $[0,w_{gripper}]$ to avoid collisions in clutter scenes, where $w_{gripper}$ represents the normalized maximal width determined by the labels (equal to the gripper width). Due to the symmetry of the parallel grippers and the robot workspace safety, $(\theta,\gamma,\beta)$ are all constrained in $[-\frac{\pi}{2},\frac{\pi}{2}]$.
IV Method
IV-A Overview
As depicted in Fig. 2, we propose a novel Region-aware Grasp Framework, comprising two core components: Normalized Grasp Space (NGS) and Regional Normalized Grasp Network (RNGNet). Establishing a dependable representation for grasp detection requires methods resilient to variations in objects and scenes. Unlike global scene-level information, regional data is relatively object- and scene-agnostic, yet it offers comprehensive insights for generating regional grasps and minimizing the impact of scene and object divergences. Hence, we aim to develop a Region-aware Grasp Framework which only inputs regional patches. The key insight of the framework is decoupling grasp detection into graspable region extraction and regional grasp generation separately. By only considering regional geometries around the grasps inside, this novel region-aware approach simplifies the grasp learning task and forces the model to learn grasp-relevant regional features. This approach not only enhances the efficiency of the method but also improves its generalizability. Further details of our method is provided in the subsequent subsections.
IV-B Normalized Grasp Space
The creation of the Normalized Grasp Space comprises three key stages: Depth-Adaptive Patch Extraction, Space Normalization, and Patch Scale Randomization. Depth-Adaptive Patch Extraction adapts the size of the receptive fields based on camera distance and object scale, ensuring consistent patch shapes while maintaining invariance to real-world distances. Space Normalization then aligns these patches to produce a representation that promotes both invariance and equivariance, enhancing the model’s generalizability and transferability. Finally, Patch Scale Randomization introduces variability in patch sizes to mimic different gripper widths and incorporates potential noise from the extraction process, thereby bolstering the model’s robustness.
IV-B1 Depth-Adaptive Patch Extraction
@@
As shown in Fig. 3(A), considering Depth-adaptive Patch Extraction, we utilize the whole RGBD image $\boldsymbol{\chi}\in\mathbb{R}^{H\times W\times 4}$ and a series of patch centers $\mathbf{p}_{i}=(x_{i},y_{i},z_{i}),i=1...N$ as input, aiming at acquiring corresponding raw patches $\boldsymbol{P}^{raw}_{i}\in\mathbb{R}^{S\times S\times 4}$ with pixel resolution $S\times S$. To leverage the rich semantic and geometric features in the patch, rather than using a fixed patch size, inspired by the scheme proposed in [16], we design a adaptive grid generation combining with the Grasp Heatmap Model to efficiently generate adaptive sampling grids with different sizes centered at the patch centers. During the process, the Grasp Heatmap Model from [3] serves to locate potential patch centers. Similar to the setting in [17], the principle is to generate adaptive receptive fields for different patches and keep the receptive fields invariant to real-world distances. Thus, the patch sizes in the original image can be formulated as below:
@@
where $r_{i}$ is the patch image side length for the $i^{th}$ patch in the original RGB image, $z_{i}$ is the depth distance from the camera frame to the $i^{th}$ patch center (shown in Fig. 3(A)(B)), and $F$ is the camera focal length along the corresponding axis. It is noteworthy that $w_{ref}$ is the reference receptive field size in the real world. To ensure the local patch size is adequate to infer grasps and avoid potential collisions, $w_{ref}$ is set to $2*w_{gripper}$. Then the regular spatial grids will be rescaled and applied to the input RGBD image, resulting in the depth-adaptive local patches. To keep the correct spatial position, the pixels in the patches are sampled using the nearest rule. No interpolation is conducted during the patch extraction process, ensuring a more consistent result when dealing with patches with different resolutions.
Figure 3: (A): Depth-Adaptive Patch Extraction demonstrations for two plaster statues of different sizes. (B): Extracted patches after adaptive grid sampling. (C): The corresponding Normalization Grasp Space. (D): The regional grasp representation as $(x,y,z,\theta,\gamma,\beta,w)$.
IV-B2 Space Normalization
@@
Accordingly, we obtain a series of raw image patches $\boldsymbol{P}_{i}^{raw}\in\mathbb{R}^{S\times S\times 4}$. During the Space Normalization, our goal is to acquire a more structural and aligned representation among patches while considering the physical size of the given gripper. Thus, firstly we convert the 4-channel raw RGBD image patches into the 6-channel RGBXYZ form $\boldsymbol{P}_{i}\in\mathbb{R}^{S\times S\times 6}$ using the camera intrinsics and imaging model, in which XYZ denotes the 3D coordinate maps (equivalent to the scene point cloud). For clarity, because RGB channels require no extra normalization, we omit the RGB channels and only consider the XYZ channels in the following discussion about Normalized Grasping Space. Then as is shown in Fig. 3(C), the normalized region patch $\boldsymbol{P}^{*}_{i}$ can be obtained with the patch centers $\mathbf{p}_{i}$ and the given gripper size $w_{gripper}$ as below:
$\boldsymbol{P}^{*}_{i}=\frac{\boldsymbol{P}_{i}-\mathbf{p}_{i}}{w_{gripper}}.$ (3)
@@
Notably, input normalization is only one-half of the construction of Normalization Grasp Space. We further consider filtering and normalizing the grasp labels inside $\boldsymbol{P}^{*}_{i}$. As mentioned in the last section, our 6-DoF grasp pose representation $\boldsymbol{g}=(x,y,z,\theta,\gamma,\beta,w)$ is equivalent to the representation in the SE(3) space $\boldsymbol{g}=(\mathbf{t},\mathbf{R},w)$, where $\mathbf{t}\in\mathbb{R}^{3\times 1}$ denotes the translation, $\mathbf{R}\in\mathbb{R}^{3\times 3}$ is the rotation matrix in the camera frame, $w$ denotes the grasp width. Thus, $\mathbf{R}$ can be represented as combination of Euler angles [1]:
$\mathbf{R}=\mathbf{R}_{z}(\theta)\mathbf{R}_{x}(\beta)\mathbf{R}_{y}(\gamma),$ (4)
where $\mathbf{R}_{z}(\theta)$ represents the rotation of $\theta$ radians around the z-axis in the camera frame, and similar notations apply for the other two rotations.
@@
Considering the reduced reliability of grasps positioned away from the center $\mathbf{p}_{i}$ , we limit our scope to grasp labels located within a spherical boundary of $0.2\times w_{gripper}$ centered at $\mathbf{p}_{i}$. Such distance boundary is consistent with the widely adopted grasp coverage criterion in [29, 38, 3], in which two grasps are regarded as equivalent if their center distance is less than a threshold. The grasps, denoted as $\boldsymbol{G}^{*}_{i}=\{\boldsymbol{g}^{*}_{j}\}$, where $\boldsymbol{G}^{*}_{i}$ is the set of normalized grasp labels $\boldsymbol{g}^{*}_{j}$:
$\boldsymbol{g}^{*}_{j}$ $=(\mathbf{t}^{*}_{j},\mathbf{R}^{*}_{j},w^{*}_{j})$
$=(\frac{\mathbf{t}_{j}-\mathbf{p}_{i}}{w_{gripper}},\mathbf{R}_{j},\frac{w_{j}}{w_{gripper}}),$ (5)
$\boldsymbol{G}^{*}_{i}$ =$=\{\boldsymbol{g}^{*}_{j}|\left\lVert\mathbf{t}^{*}_{j}\right\rVert<0.2\}.$ (6)
As Fig. 3(D) indicates, our region-aware framework only focuses on the grasps near the patch centers, which makes the problem statement more clear and more concentrated.
@@
Then for the regional 6-DoF grasp detection problem, the defined grasp function is $f:\boldsymbol{P}^{*}_{i}\to\boldsymbol{G}^{*}_{i}$. Our goal is learn a network $\hat{f}_{\lambda}:\boldsymbol{P}^{*}_{i}\to\hat{\boldsymbol{G}}^{*}_{i}$ with parameters $\lambda$ to fit $f$. By inverting the normalizing process, we can finally get the grasp predictions $\hat{\boldsymbol{G}}_{i}$ for this patch in the camera frame.
Figure 4: Visualization of the characteristics of Normalized Grasp Space.
IV-B3 Characteristics of Normalized Grasp Space
@@
As shown in Fig. 4, constraining the grasp detection problem in the Normalized Grasp Space brings beneficial characteristics for the 6-DoF grasp detection within the regional patches task as follows. For clarity, we omit the patch index $i$ and denotes the grasp labels within the transformed patches as $\boldsymbol{G}^{*}_{t}$:
@@
  • Translation-invariance: When the patch $\boldsymbol{P}_{i}$ translates with a range $\Delta\mathbf{t}$, benefiting from the normalization process, we get the same normalized patch $\boldsymbol{P}^{*}_{i}$ and the grasps inside the patch keep invariant:
    $\boldsymbol{G}^{*}_{t}=f(\boldsymbol{P}^{*}_{t})$ $=f(Norm(\boldsymbol{P}+\Delta\mathbf{t}))$
    $=f(Norm(\boldsymbol{P}))$
    $=f(\boldsymbol{P}^{*})=\boldsymbol{G}^{*}.$ (7)
    It is noteworthy that the grasp translation-invariance within the patches can be seamlessly integrated with other equivariant patch location models (heatmap, segmentation, tracking and …) to achieve grasp detection translation-equivariance at the scene level.

### v2

# Region-aware Grasp Framework with Normalized Grasp Space for Efficient 6-DoF Grasping

@@ Abstract
A series of region-based methods succeed in extracting regional features and enhancing grasp detection quality. However, faced with a cluttered scene with potential collision, the definition of the grasp-relevant region stays inconsistent, and the relationship between grasps and regional spaces remains incompletely investigated. In this paper, we propose Normalized Grasp Space (NGS) from a novel region-aware viewpoint, unifying the grasp representation within a normalized regional space and benefiting the generalizability of methods. Leveraging the NGS, we find that CNNs are underestimated for 3D feature extraction and 6-DoF grasp detection in clutter scenes and build a highly efficient Region-aware Normalized Grasp Network (RNGNet). Experiments on the public benchmark show that our method achieves significant $>20\%$ performance gains while attaining a real-time inference speed of approximately 50 FPS. Real-world cluttered scene clearance experiments underscore the effectiveness of our method. Further, human-to-robot handover and dynamic object grasping experiments demonstrate the potential of our proposed method for closed-loop grasping in dynamic scenarios.
Keywords: 6-DoF Grasping, RGBD Perception, Normalized Space, Heatmap
1 Introduction
@@
Grasping, a fundamental task for robotic manipulation, has seen significant advancements. Notably, recent strides in deep-learning methodologies have paved the way for data-driven grasp detection techniques. These techniques, capable of directly deducing the positions and rotations of available grasps, offer superior generalization to novel scenes compared to the traditional template-based methods. Meanwhile, 6-DoF grasping has grown significantly because of its effectiveness in applications and scenarios that require accurate and reasonable manipulation of objects.
@@
Towards efficient 6-DoF grasping, most grasp detection methods treat the task as a detection-style problem involving the acquisition of optimal grasp positions and rotations from a single-view RGBD image or point cloud. Most of them process the global scene information to generate all possible grasps in the scene. Typical works like [1, 2] extract scene geometric features and regress grasps directly. The crucial problem is the redundancy and ambiguity of the entire scene information, which is time-consuming. Facing the challenge of grasping in cluttered scenes, recent methods [3, 4, 5, 6, 7, 8] encode global per-points features to infer potential grasp centers and segment graspable regions by integrating the regional points with the global features. While global-to-local schemes have demonstrated effectiveness, leveraging regional information without a consistent representation may undermine the generalization capability for novel scenes. The primary challenge lies in devising a method to extract regional information in a consistent representation.
@@
In this work, we propose a novel Region-aware Grasp Framework decoupling grasp detection into graspable region extraction and follow-up regional grasp generation. Specifically, we propose Normalized Grasp Space (NGS), a unified representation bridging the patch extraction and regional grasp detection. With extracted region information, the grasp detection problem within NGS can be formulated into a normalized form, which shows the invariance to the gripper size and the robustness to perturbations on the grasp center localization. Thus, considering grasp detection from a region-aware viewpoint, we propose Regional Normalized Grasp Network (RNGNet) to extract grasp-relevant geometric features and generate rotation heatmaps, demonstrating enhanced robustness and efficiency. Quantitative experiment results show a significant over 20% improvement in the grasp Average Precision (AP) on the public benchmark with 10×10$\times$10 × faster inference speed than the former SOTA, AnyGrasp [9]. Real-world clutter clearance experiments validate our framework’s robustness and generalizability. Further demonstration of grasping in dynamic scenes proves the efficiency and flexibility of our proposed region-aware framework and the potential for more complex tasks like target-oriented manipulation. The contributions of the paper mainly include:
@@
1) A novel Region-aware Grasp Framework for 6-DoF grasp detection in cluttered scenes, decoupling grasp detection tasks into two-stage heatmap prediction and enhancing method efficiency.
@@
2) Normalized Grasp Space (NGS), a novel unified space defined for general parallel gripper grasp detection, empowering aligned and scale-invariant grasp detection in regions.
@@
3) A highly efficient Region Normalized Grasp Network (RNGNet) aiming at extracting region-aware features and predicting high-quality 6-DoF grasps via rotation heatmap prediction.
2 Related Work
@@
Vision-based 6-DoF Grasp Detection: Research has focused on detecting 6-DoF grasps primarily relying on single-view RGBD images or point clouds. Pioneer works ten Pas et al. [10], Liang et al. [11] mainly follow a sample-evaluation strategy, generating numerous grasp proposals and selecting optimal grasps through an evaluation model. Qin et al. [2] employ PointNet++ [12] for per-point feature extraction and direct regression of grasp parameters. Recent advancements, benefiting from large-scale datasets [3, 13, 14], adopt an end-to-end strategy. Wang et al. [6] introduces point-wise graspness to represent grasp location probabilities and aggregate regional features for rotation prediction. HGGD [15] employs 2D CNNs to generate heatmaps for grasp locations and integrate 2D semantics with 3D geometric features to configure 6D grasps. Diverging from existing approaches, our method employs a region-aware scheme forcing the network to capture scene-independent regional features, enhancing efficiency and generalization capability.
Region Normalization for Grasp Detection: Traditional 4-DoF grasp detection approaches [16, 17] conduct image-based normalization or crop the image patches with the exact sizes [18] to improve model performance. However, such 2D frameworks are limited, overlooking critical 3D geometric information and sensitive to camera viewpoints. For 6-DoF grasp detection, the utilization of regional 3D geometry is also not unique. However, most methods operate under a fixed gripper setting and extract local regions of predetermined sizes [2, 3, 6, 5]. These methods struggle to generalize or require model retraining when applied to grippers with different sizes or novel scenes. We normalize regions and corresponding grasps according to adjustable gripper sizes to obtain consistent 3D normalized grasp spaces, allowing more efficient grasp detection and adaptation for challenging applications, such as dynamic grasping.
@@
Figure 1: Region-aware Grasp Framework. With an RGBD image input, Grasp Heatmap Model (GHM) [15] predicts grasp location heatmap. Subsequently, graspable patches are extracted through an adaptive grid generator, which samples mesh grids based on the location heatmap and the gripper scale. Local patches are then converted into Normalized Grasp Space and fed into the Regional Normalized Grasp Network to predict regional rotation heatmaps and grasps. Finally, we apply the inverse of the normalization process and obtain the scene-level grasps.
3 Method
@@
Similar to prior works [15, 19, 3], we focus on the problem of 6-DoF grasp detection of parallel grippers in clutter scenes from a single-view RGBD image $\boldsymbol{\chi}\in\mathbb{R}^{H\times W\times 4}$ as input. Inspired by [15], we adopt the grasp representation as $\boldsymbol{g}=(x,y,z,\theta,\gamma,\beta,w)$, in which $(x,y,z)$ is the 3-DoF translations of grasp centers, $w$ is the grasp width and $(\theta,\gamma,\beta)$ compose the 3-DoF intrinsic rotation as Euler angles. $w$ is predicted within the range of $[0,w_{gripper}]$ to avoid collisions in clutter scenes. $(\theta,\gamma,\beta)$ are all constrained in $[-\frac{\pi}{2},\frac{\pi}{2}]$.
3.1 Normalized Grasp Space
Figure 2: Depth-Adaptive Patch Extraction and grasp representation.
@@ 3.1 Normalized Grasp Space
Depth-Adaptive Patch Extraction: As shown in Fig. 2(A)(B), we utilize the RGBD image $\boldsymbol{\chi}\in\mathbb{R}^{H\times W\times 4}$ and patch centers $\mathbf{p}_{i}=(x_{i},y_{i},z_{i})$ sampled from grasp location heatmap predicted by [15] as input, aiming at acquiring corresponding patches $\boldsymbol{P}^{raw}_{i}\in\mathbb{R}^{S\times S\times 4}$. To better leverage geometric features and avoid camera viewpoint variations, rather than using a fixed patch size, we design an adaptive grid generator that efficiently generates sampling grids with different sizes centered at the centers. Inspired by the setting in [20], the principle is to generate adaptive receptive fields for different patches and keep the receptive fields invariant to real-world distances. Thus, the patch sizes can be formulated as:
@@
where $r_{i}$ is the patch image side length for the $i^{th}$ patch, $z_{i}$ is the depth to the $i^{th}$ center in the camera frame, and $F$ is the camera focal length. It is noteworthy that $w_{ref}$ is the reference receptive field size in the real world. To ensure the local patch size is adequate to infer grasps and avoid potential collisions, $w_{ref}$ is set to $2*w_{gripper}$. Then the regular spatial grids will be rescaled and applied to the input RGBD image, resulting in the depth-adaptive local patches.
@@
Space Normalization: Normalized Space has been proved helpful for neural network learning. During Space Normalization, our goal is to acquire a more structural and aligned representation among patches and regional grasps. Thus, firstly we convert the 4-channel raw RGBD image patches into the 6-channel RGBXYZ form $\boldsymbol{P}_{i}\in\mathbb{R}^{S\times S\times 6}$ using the camera intrinsics and imaging model, in which XYZ denotes the 3D coordinate maps. For clarity, because RGB channels require no extra normalization, we omit the RGB channels and only consider the XYZ channels in the following discussion. Then as is shown in Fig. 2(C), the normalized region patch $\boldsymbol{P}^{*}_{i}$ can be obtained with the patch centers $\mathbf{p}_{i}$ and the given reference receptive field $w_{ref}$ as below:
$\boldsymbol{P}^{*}_{i}=\frac{\boldsymbol{P}_{i}-\mathbf{p}_{i}}{w_{ref}}.$ (2)
@@
Notably, input normalization is only part of the Normalization Grasp Space. We filter and normalize the grasp labels inside $\boldsymbol{P}^{*}_{i}$. Our 6-DoF grasp pose representation $\boldsymbol{g}=(x,y,z,\theta,\gamma,\beta,w)$ is equivalent to the representation in the SE(3) space $\boldsymbol{g}=(\mathbf{t},\mathbf{R},w)$, where $\mathbf{t}\in\mathbb{R}^{3\times 1}$ denotes the translation, $\mathbf{R}\in\mathbb{R}^{3\times 3}$ is the rotation matrix in the camera frame, $w$ denotes the grasp width. Thus, $\mathbf{R}$ can be represented as Euler angles: $\mathbf{R}=\mathbf{R}_{z}(\theta)\mathbf{R}_{x}(\beta)\mathbf{R}_{y}(\gamma)$, where $\mathbf{R}_{z}(\theta)$ represents the rotation of $\theta$ radians around the z-axis and similar notations apply for the other two rotations.
@@
We limit our scope to grasp labels located within a spherical boundary of $0.1*w_{ref}$ centered at $\mathbf{p}_{i}$. Such distance boundary is consistent with the widely adopted grasp coverage criterion in [21, 4, 15], in which two grasps are regarded as equivalent if their center distance is less than a threshold. The grasps, denoted as $\boldsymbol{G}^{*}_{i}=\{\boldsymbol{g}^{*}_{j}\}$, where $\boldsymbol{G}^{*}_{i}$ is the set of normalized grasp labels $\boldsymbol{g}^{*}_{j}$:
$\boldsymbol{g}^{*}_{j}=(\mathbf{t}^{*}_{j},\mathbf{R}^{*}_{j},w^{*}_{j})=(\frac{\mathbf{t}_{j}-\mathbf{p}_{i}}{w_{ref}},\mathbf{R}_{j},\frac{w_{j}}{w_{ref}}),$
$\boldsymbol{G}^{*}_{i}=\{\boldsymbol{g}^{*}_{j}|\left\lVert\mathbf{t}^{*}_{j}\right\rVert<0.1\}.$ (3)
@@
As Fig. 2(D) indicates, our region-aware framework only focuses on the grasps near the patch centers, making the problem statement clearer and more concentrated. Then, for the regional 6-DoF grasp detection problem, the defined grasp function is $f:\boldsymbol{P}^{*}_{i}\to\boldsymbol{G}^{*}_{i}$. Our goal is learn a network $\hat{f}_{\lambda}:\boldsymbol{P}^{*}_{i}\to\hat{\boldsymbol{G}}^{*}_{i}$ with parameters $\lambda$ to fit $f$. By inverting the normalizing process, we can finally get the grasp predictions $\hat{\boldsymbol{G}}_{i}$ for this patch in the camera frame.
@@
Characteristics of Normalized Grasp Space: As shown in Fig. 3, constraining the grasp detection problem in the Normalized Grasp Space brings beneficial characteristics as follows. For clarity, we omit the patch index $i$ and denotes the grasp labels within the transformed patches as $\boldsymbol{G}^{*}_{\mathcal{T}}$.
Figure 3: Visualization for the characteristics of Normalized Grasp Space.
@@
1) Translation-invariance: When the patch $\boldsymbol{P}_{i}$ translates with $\Delta\mathbf{t}$, we get the same normalized patch $\boldsymbol{P}^{*}_{i}$ and the grasps inside the patch keep invariant:
$\boldsymbol{G}^{*}_{\mathcal{T}}=f(\boldsymbol{P}^{*}_{\mathcal{T}})=f(\text{Norm}(\boldsymbol{P}+\Delta\mathbf{t}))=f(\text{Norm}(\boldsymbol{P}))=f(\boldsymbol{P}^{*})=\boldsymbol{G}^{*}.$ (4)